OpenStudy (anonymous):

0.5b + 4 = 2(b+2) help please

OpenStudy (anonymous):

2( b/4 +2) = 2(b+2) b/4 + 2 = b + 2 0 = b-b/4 0 = 3b/4 b=0

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Welcome..